You are interested in purchasing a wide screen television at Target. The radio width of the screen to the height of the screen is 16 to 9. If a certain model you are considering has a screen width of 48 inches, what would be the height of this screen.
Answer by Alan3354(69443)

You are interested in purchasing a wide screen television at Target. The radio width of the screen to the height of the screen is 16 to 9. If a certain model you are considering has a screen width of 48 inches, what would be the height of this screen.
-------------
w/h = 16/9
---
48/h = 16/9
h = 27 inches